jigglypuff314 (jigglypuff314):

One pound of cashews and 2 pounds of peanuts are added to 3 pounds of a store-bought cashew/peanut mix. If the total is a 6-pound mix that is 2/5 cashews by weight, what fraction of the store-bought mix were cashews?

OpenStudy (lovelyanna):

Woah.. confusing :/

OpenStudy (lovelyanna):

Emily.. I don't know >:'( ....

jigglypuff314 (jigglypuff314):

that's alright ;) someone will figure it out...

OpenStudy (anonymous):

Okay, @jigglypuff314, have to agree with @LovelyAnna that the question is phrased a little oddly. However, I think this may be what you are looking for: \[\frac{ 1 }{ 3 }+x=\frac{ 2 }{ 5 }\] Solve for x and I get:\[x=\frac{ 1 }{ 15 }\]

OpenStudy (lovelyanna):

So sorry.. :(

jigglypuff314 (jigglypuff314):

@HomeschoolGrad that's what I got too but the answer key had gotten 7/15 ... do you have an idea of how they would have gotten that?

OpenStudy (anonymous):

:S could it be a mistake in the answer key? You would be surprised how many times I've driven myself crazy just because of a mistake in an answer key. Like I said though, the question is phrased a little weirdly. Maybe there is something we're not understanding. :(

jigglypuff314 (jigglypuff314):

alright :) I'll ask my teacher then :) thanks for helping! :D

OpenStudy (lovelyanna):

Sorry! :P Good luck!

jigglypuff314 (jigglypuff314):

It's alright ;)

OpenStudy (lovelyanna):

So, how was your Saturday?

jigglypuff314 (jigglypuff314):

spent my saturday playing flute for a craft fair ;) it was fun :) but playing flute for 3 hours really hurts my mouth :P

OpenStudy (lovelyanna):

Cool :) Lol, be careful! ;)

jigglypuff314 (jigglypuff314):

yeah ;) I have an audition next saturday :P

OpenStudy (lovelyanna):

Awesome! Good luck and do your best! :D

jigglypuff314 (jigglypuff314):

thanks :)

OpenStudy (lovelyanna):

I was sick yesterday :/ Now I'm in a hurry to finish my History.. I couldn't do it yesterday.. was too tired

jigglypuff314 (jigglypuff314):

I suppose I can try to help :P but my chances of getting stuff correct for history questions is 50% ... just saying ;)

OpenStudy (lovelyanna):

Lol, thanks :) Someone is helping right now, you can come join. Let me tag you! :)

hartnn (hartnn):

its indeed 7/15

hartnn (hartnn):

' 6-pound mix that is 2/5 cashews by weight' means 6*2/5 = 12/5 pounds cashew in total. 1 pound was added to the mix, so, 12/5-1 = 7/5 pound of cashew was already present in that 3 pound mix 7/5 pound in 3pounds gives the ratio of (7/5)/3 = 7/15